ANALYSIS OF THE C^{2}Π − X^{2}Σ^{+} (0,0) BAND OF AlO.

\underline{JONATHAN P. TOWLE}, Smithsonian Astrophysical Observatory, 60 Garden Street MS 72, Cambridge, MA 02138, USA.

In recent years the C^{2}Π state of AlO has been the subject of some controversy. In 1994 Towle et al. ^{a} published the results and analysis of a high resolution study of the C^{2}Π − X^{2}Σ^{+} (0,0) band. Later in 1994 Towle et al. ^{b} published a reanalysis of that data in combination more data from Singh and Sakena ^{c} The reanalysis yielded molecular constants for the C^{2}Π state that were in agreement with those published in ^{a} to within experimental accuracy. Ito et al. ^{d} have reanalyzed the data from ^{c} and published a set of molecular constants that are markedly different from those of Towle et al. ^{a} These differences will be discussed.

aJ.P. Towle, A.M. James, O.L. Bourne and B. Simard, J. Mol. Spectrosc. 163 (1994) 300.

bJ.P. Towle, A.M. James, O.L. Bourne, B. Simard, M. Singh and M.D. Sakena, J. Mol. Spectrosc. 167 (1994) 472.

cM. Singh and M.D. Sakena, Can. J. Phys. 61 (1983) 1347.

dN. Ito, H. Ito and K. Kuchitsu, Chem. Phys. Lett., 276 (1997) 139.